# CoLA: Compute-Efficient Pre-Training of LLMs via Low-Rank Activation

### v2

@@ 3.3 Computing Efficiency
Table 3: Estimated compute of a single LLaMA decoder layer for different pre-training methods. Results combine forward, backward and any additional compute occurred at optimizer step.
We analyze and compare the computational complexity of CoLA with other efficient pre-training methods based on the LLaMA architecture. We adopt a similar notion from Kaplan et al. (2020), where a general matrix multiply (GEMM) between an $M\times N$ matrix and an $N\times K$ matrix involves roughly $2MNK$ add-multiply operations. We denote the model inner width as $d$, and the inner width of the feed-forward layer as $d_{\text{ff}}$. For simplicity, we only show non-embedding calculations of a single sequence with token batch size of $n$ for each decoder layer. This is because the total computation scales only linearly with the number of layers $n_{\text{layer}}$ and the number of sequences $n_{\text{seq}}$. Furthermore, lower-order cheap operations of complexity $\mathcal{O}(nd)$ or $\mathcal{O}(nd_{\text{ff}})$ are omitted, such as bias, layer norm, non-linear function, residual connection, and element-wise product.
We show the detailed cost of the full-rank training in Table. 2. Notice that we apply the 2×2$\times$2 × rule when calculating the backward cost. This is because for each forward GEMM that Eq. (2) describes, two GEMMs are needed to compute gradients for both the weight matrix $\mathbf{W}$ and the input $\mathbf{x}$, and are of the same cost the forward GEMM, i.e.,
@@ 4.1 Memory Breakdown in Pre-Training
We assume a common notion that training modern transformers with Adam (or AdamW) involves four key memory components Zhao et al. (2024); Han et al. (2024): model parameters (1×1$\times$1 ×), gradients (1×1$\times$1 ×), optimizer states (2×2$\times$2 ×), and activations (14×1$\sim$ 4$\times$1 ∼ 4 ×). We focus on the scenario where the memory cost determined by the model size is not on the extreme limit of the GPU. We argue that this is rather realistic, since the model size and the minimum required tokens should scale up simultaneously during pre-training Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024). A tiny batch size on a single GPU would be impractical. Therefore, we analyze memory usage on a 40-GB A100 or a 94-GB H100 GPU with a fairly large sequence batch size. Fig. 5 & 6 show that activations dominate memory usage in this setup.
@@ 4.2 CoLA Enables Efficient Checkpointing
Methods Memory Re-Compute
Full-Rank $20nd+2n^{2}h$ N/A
Vanilla GCP $nd$ $23nd^{2}+4n^{2}d$
CoLA $17.5nd+2n^{2}h+14nr$ N/A
CoLA-M $2nd+7nr$ $18.5ndr+4n^{2}d$
Figure 7: We show how memory reduction scales with the re-computation in full-rank training with GCP and compare with CoLA-M. With similar gains on memory efficiency, CoLA-M effectively reduces re-compute by 4.6×4.6$\times$4.6 ×, enabling compute efficient checkpointing.
@@
Gradient checkpointing (GCP) Chen et al. (2016) is a system-level technique that reduces memory usage by selectively storing (“checkpointing”) only a subset of intermediate activations during the forward pass. When the backward pass begins, the missing activations are recomputed on the fly instead of being stored in memory, thereby lowering the memory cost. A vanilla (also the most effective) implementation of GCP in LLM pre-training is to save merely the input and output of each transformer block, and re-compute everything within each block during the backward step. Some works have investigated the optimal selection of checkpoints through both empirical and compiler view Feng and Huang (2021); He and Yu (2023). Such techniques can also be developed for CoLA, and are beyond the scope of this paper.
Motivated by the bottleneck structure of CoLA, we implement CoLA-M as saving only the low-rank activations (red circles in Fig. 4), and re-compute the up projections, and (if applicable) the self-attention (painted in sketch in Fig. 4) during the backward pass. This reduces the re-computation cost to half of the CoLA forward. We analyze the memory and re-computation cost using the same notions as in Section 3.3 and denote $h$ as the number of attention heads. We further simplify the analysis under LLaMA architecture by uniformly assuming $d_{\text{ff}}\approx 2.5d$. The memory and re-computation overhead are shown in Table 4. We refer the detailed analysis to Appendix C.
Although delicate optimizations of GCP is beyond our scope, we show in Fig. 7 the quantitative results and scaling behavior of GCP on LLaMA-1B when applying a heuristic checkpointing strategy. We observe that CoLA-M greatly reduces re-computation cost by $\bf 4.6\boldsymbol{\times}$ while achieving similar memory saving (18.94GB) as vanilla GCP (20.25GB).
@@ 5 Experiments
Full-Rank 34.06 1×1$\times$1 × 24.36 1×1$\times$1 × 18.80 1×1$\times$1 ×
Control 37.73 0.4×0.4$\times$0.4 × 27.05 0.5×0.5$\times$0.5 × 20.53 0.4×0.4$\times$0.4 ×
CoLA 34.04 0.4×0.4$\times$0.4 × 24.48 0.5×0.5$\times$0.5 × 19.40 0.4×0.4$\times$0.4 ×
31.52 0.7×0.7$\times$0.7 × 23.97 0.7×0.7$\times$0.7 × 18.32 0.7×0.7$\times$0.7 ×
@@ 5.1 Pre-Training within Compute-Optimal
We implement CoLA and CoLA-M by parameterizing all MLP layers and all projection layers in attention with auto-encoders [i.e. Eq. (3)], and keep all other parameters and operations unchanged. We use AdamW optimizer and cosine annealing learning rate scheduler Loshchilov and Hutter (2016) with warm-up. We refer detailed configurations to Appendix D.
Table 5 compares our methods and other efficient pre-training techniques in terms of validation perplexity, parameter size, and estimated memory usage of model, gradients and optimizer states. CoLA has the smallest model size, thereby consumes the least memory, and performs on-par with full-rank baselines. CoLA uniformly surpasses other efficient training baselines in both efficiency and performance. Table 6 compares the validation perplexity on the 7B model for 150k steps [^3] Due to resources constraints, 7B models are trained below compute optimal budget Zhao et al. (2024); Han et al. (2024).. CoLA(-M) significantly outperforms 8-bit Adam/GaLore by 12.73 vs ~14.6, while saving two-third memory.
Scaling Behavior: Table 7 shows how CoLA might be improved when compute is scaled up. The default rank choices reduce half the computing cost, without harming the model performance. Meanwhile, if we relax the computing restriction and moderately increase the rank, then CoLA outperforms full-rank training in all three scales, while still being fairly smaller and reducing the computing cost. One might argue that full-rank training can also be scaled down to a similar computing cost of CoLA and might perform similarly. We implement such baselines in Table 7 and refer this setup to “Control". We typically reduce the number of layers or the model width of full-rank models to scale down their computing cost. We find empirically that they reduce performance significantly and dramatically underperform CoLA.
@@ 5.2 Pre-Training beyond Compute-Optimal
LLaMA-350M with 51B Tokens: We prolong the training duration by 8$\times$ of the compute-optimal budget for both CoLA [^4] We choose CoLA at 0.7×0.7$\times$0.7 × compute of full-rank baseline, as its superior performance observed in Table 7. and full-rank LLaMA at 350M scale. This results in 51B total training tokens. CoLA continues outperforming full-rank baseline on validation perplexity of 13.96 vs 14.47, consistent with results at compute-optimal observed from Table 7.
@@
These results further demonstrate CoLA’s effectiveness across both encoder/decoder architectures, both compute-optimal/over-train settings, and different activations (GeLU, Swish and SwiGLU).
@@ 5.3 Training/Inference System Performance
Full-Rank 69.84 12,365 1×1$\times$1 × 84.94 5,810 1×1$\times$1 ×
Vanilla GCP 14.89 8,799 1.68×1.68$\times$1.68 × 52.49 4,357 1.67×1.67$\times$1.67 ×
@@
Not just for training, CoLA also speeds up inference and reduces memory cost. Table 11 (Appendix E.2) shows that CoLA off-the-shelf improves inference throughput by up to $\bf 1.64\boldsymbol{\times}$ while reducing memory cost by up to $\bf 1.67\boldsymbol{\times}$.
@@ 6 Conclusions
We have proposed CoLA, and its memory efficient variant CoLA-M, to achieve collectively parameter, computing and memory efficiency at both training and inference time for large foundation models. CoLA effectively reduces $\bf 2\boldsymbol{\times}$ model size and computing cost while preserving full-rank level performance. CoLA-M trades minimum overhead for state-of-the-art memory reduction, while still improving training throughput over full-rank baselines. Crucially, CoLA is promising to save substantial GPU resources in LLM industry. This work has been focused on dense architectures. In the future, it is worth extending CoLA to the mixture-of-expert (MoE) architecture.
@@ Appendix A Observation of Low-Rank Activation in Pre-Trained GPT2
In this section, we further show the low-rank structure in model activations evaluated on a pre-trained GPT-2 Radford et al. (2019) small. The evaluation is conducted with sequence batch size of 64 and sequence length of 1024. We fix $\alpha=0.95$ throughout this section. Similar patterns are observed from the attention layers (Fig. 91011). The low-rank nature of activations is evident across all the different components of the model. This suggests that despite the high-dimensional representations, the effective dimensionality of the activations remains constrained.
Figure 9: Activation Spectrum of Attention Layer (Q)
Figure 10: Activation Spectrum of Attention Layer (K)
Figure 11: Activation Spectrum of Attention Layer (V)
@@ Appendix B Detailed Compute Analysis
where $\oplus$ denotes the scatter-add operator, $\mathcal{I}$ and $\mathcal{V}$ denote the indices and values of non-zero elements in $\mathbf{S}$. This implementation avoids instantiating a full sized $\mathbf{S}$, instead keeping only the non-zero elements. However, this introduces non-trivial reconstruction cost of $\mathbf{BA}$ in every step. And if we further denote $\tilde{\mathbf{W}}=\mathbf{BA}\oplus_{\mathcal{I}}\mathcal{V}$, then the forward data-flow that starts from $\tilde{\mathbf{W}}$ is the same as in the full-rank training, as well as the backward data-flow that ends at $\tilde{\mathbf{W}}$. Therefore, the total compute of SLTrain should be $C_{\text{full-rank}}$ plus reconstructing $\tilde{\mathbf{W}}$, and its corresponding 2×2$\times$2 × compute during backward, i.e.,
@@ Appendix C Detailed Memory Analysis
We continue using the notions defined in Section. 4.2 and start with the activation memory of full-rank training:
@@
Appendix D Hyper-Parameters
@@ E.1 Ablation Study
CoLA w/ Only Low-Rank $\sigma$ – Reduced 35.41 25.90 20.50

### v3

@@ 3.4 Computing Efficiency
Table 3: Estimated computing cost of a single LLaMA decoder layer. Results combine forward, backward and any additional compute occurred at optimizer step.
We analyze and compare the computational complexity of CoLA with other pre-training methods based on the LLaMA architecture. We adopt a similar notion from Kaplan et al. (2020), where a general matrix multiply (GEMM) between an $M\times N$ matrix and an $N\times K$ matrix involves roughly $2MNK$ add-multiply operations. We denote the model inner width as $d$, and the inner width of the feed-forward layer as $d_{\text{ff}}$. For simplicity, we only show non-embedding calculations of a single sequence with token batch size of $n$ for each decoder layer. This is because the total computation scales only linearly with the number of layers $n_{\text{layer}}$ and the number of sequences $n_{\text{seq}}$. Furthermore, lower-order cheap operations of complexity $\mathcal{O}(nd)$ or $\mathcal{O}(nd_{\text{ff}})$ are omitted, such as bias, layer norm, non-linear function, residual connection, and element-wise product.
We show the detailed cost of the full-rank training in Table. 2. Notice that we apply the $2\times$ rule when calculating the backward cost. This is because for each forward GEMM that Eq. (2) describes, two GEMMs are needed to compute gradients for both the weight matrix $\mathbf{W}$ and the input $\mathbf{x}$, and are of the same cost the forward GEMM, i.e.,
@@ 4.1 Memory Breakdown in Pre-Training
We assume a common notion that training modern transformers with Adam (or AdamW) involves four key memory components Zhao et al. (2024); Han et al. (2024): model parameters ($1\times$), gradients ($1\times$), optimizer states ($2\times$), and activations ($1\sim 4\times$). We focus on the scenario where the memory cost determined by the model size is not on the extreme limit of the GPU. We argue that this is rather realistic, since the model size and the minimum required tokens should scale up simultaneously during pre-training Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024). A tiny batch size on a single GPU would be impractical. Therefore, we analyze memory usage on a 40-GB A100 or a 94-GB H100 GPU with a fairly large sequence batch size. Fig. 5 & 6 show that activations dominate memory usage.
@@ 4.2 CoLA Enables Efficient Checkpointing
Figure 7: We show how memory reduction scales with the re-computation in full-rank training with GCP and compare with CoLA-M. With similar gains on memory efficiency, CoLA-M effectively reduces re-compute by $4.6\times$, enabling compute efficient checkpointing.
@@
Gradient checkpointing (GCP) Chen et al. (2016) is a system-level technique that reduces memory usage by selectively storing (“checkpointing”) only a subset of intermediate results during the forward pass. When the backward pass begins, the missing activations are recomputed on the fly instead of being stored in memory, thereby lowering the memory cost. A vanilla (also the most effective) implementation of GCP in LLM pre-training is to save merely the input and output of each transformer block, and re-compute everything within each block during the backward step. Some works have investigated the optimal selection of checkpoints through both empirical and compiler view Feng and Huang (2021); He and Yu (2023). Such techniques can also be developed for CoLA, and are beyond the scope of this paper.
Motivated by the bottleneck structure of CoLA, we implement CoLA-M as saving only the low-rank activations (red circles in Fig. 4), and re-compute the up projections, and (if applicable) the self-attention (painted in sketch in Fig. 4) during the backward pass. This reduces the re-computation cost to half of the CoLA forward. We refer the detailed analysis to Appendix C.
Although delicate optimizations of GCP is beyond our scope, we show in Fig. 7 the quantitative results and scaling behavior of GCP on LLaMA-1B when applying a heuristic checkpointing strategy. CoLA-M greatly reduces the re-computation cost by $\bf 4.6\boldsymbol{\times}$ while achieving similar memory saving (18.94GB) as vanilla GCP (20.25GB).
@@ 5 Experiments
Full-Rank 34.06 $1\times$ 24.36 $1\times$ 18.80 $1\times$
Control 37.73 $0.4\times$ 27.05 $0.5\times$ 20.53 $0.4\times$
CoLA 34.04 $0.4\times$ 24.48 $0.5\times$ 19.40 $0.4\times$
31.52 $0.7\times$ 23.97 $0.7\times$ 18.32 $0.7\times$
@@ 5.1 Pre-Training within Compute-Optimal
We implement CoLA and CoLA-M by parameterizing all MLP layers and all projection layers in attention with auto-encoders [i.e. Eq. (3)], and keep all other parameters and operations unchanged. We use AdamW optimizer and cosine annealing learning rate scheduler Loshchilov and Hutter (2016) with warm-up. We show details to Appendix D.1.
Table 4 compares our methods and other efficient pre-training techniques in terms of validation perplexity, parameter size, and estimated memory usage of model, gradients and optimizer states. CoLA has the smallest model size, thereby consumes the least memory, and performs on-par with full-rank baselines. CoLA uniformly surpasses other efficient training baselines in both efficiency and accuracy. Table 5 compares the validation perplexity on the 7B model for 150k steps [^5] Due to resources constraints, 7B models are trained below compute optimal budget Zhao et al. (2024); Han et al. (2024).. CoLA(-M) significantly outperforms 8-bit Adam/GaLore by 12.73 vs ~14.6, while saving two-third memory.
Scaling Behavior: Table 6 shows how CoLA might be improved when compute is scaled up. The default rank choices reduce half the computing cost, without harming the model performance. Meanwhile, if we relax the computing restriction and moderately increase the rank, then CoLA outperforms full-rank training in all three scales, while still being fairly smaller and reducing the computing cost. One might argue that full-rank training can also be scaled down to a similar computing cost of CoLA and might perform similarly. We implement such baselines in Table 6 and refer this setup to “Control". We typically reduce the number of layers or the model width of full-rank models to scale down their computing cost. We find empirically that they increase perplexity (PPL) significantly and dramatically underperform CoLA.
@@ 5.2 Pre-Training beyond Compute-Optimal
[…]
These results further demonstrate CoLA’s effectiveness across both encoder/decoder architectures, both compute-optimal/over-train settings, and different activations (GeLU and Swish).
@@ 5.3 Training/Inference System Performance
Full-Rank 69.84 12,365 $1\times$ 84.94 5,810 $1\times$
Vanilla GCP 14.89 8,799 $1.68\times$ 52.49 4,357 $1.67\times$
@@
CoLA also speeds up inference and reduces memory cost. Table 12 (Appendix E.2) shows that CoLA off-the-shelf reduces inference latency and memory cost by up to $\bf 1.64\boldsymbol{\times}$ and $\bf 1.67\boldsymbol{\times}$, respectively.
@@ 6 Conclusions
We have proposed CoLA, and its memory efficient variant CoLA-M, to achieve collectively parameter, computing and memory efficiency in both pre-training and inference for large foundation models. CoLA has reduced $\bf 2\boldsymbol{\times}$ model size and computing cost while preserving full-rank level performance. CoLA-M trades minimum overhead for state-of-the-art memory reduction, while still improving training throughput over full-rank baselines. CoLA is promising to save substantial GPU resources in LLM industry. This work has focused on dense architectures. In the future, it is worth extending CoLA to the mixture-of-expert (MoE) architecture.
@@ Appendix A Observation of Low-Rank Activation in Pre-Trained GPT2
In this section, we further show the low-rank structure in model activations evaluated on a pre-trained GPT-2 Radford et al. (2019) small. The evaluation is conducted WikiText2 dataset with sequence length 1024. We fix $\alpha=0.95$ throughout this section. Similar patterns are observed from the attention layers (Fig. 91012). The low-rank nature of activations is evident across all the different components of the model. This suggests that despite the high-dimensional representations, the effective dimensionality of the activations remains constrained.
Figure 9: Spectrum of attention layer (query) output (i.e., $\mathbf{W_{q}x}$).
Figure 10: Spectrum of attention layer (key) output (i.e., $\mathbf{W_{k}x}$).
Figure 11: Spectrum of attention layer (value) output (i.e., $\mathbf{W_{v}x}$).
Figure 12: Spectrum of MLP block output (i.e., $\mathbf{W_{2}\sigma(W_{1}x)}$).
@@ Appendix B Detailed Compute Analysis
where $\oplus$ denotes the scatter-add operator, $\mathcal{I}$ and $\mathcal{V}$ denote the indices and values of non-zero elements in $\mathbf{S}$. This implementation avoids instantiating a full sized $\mathbf{S}$, instead keeping only the non-zero elements. However, this introduces non-trivial reconstruction cost of $\mathbf{BA}$ in every step. And if we further denote $\tilde{\mathbf{W}}=\mathbf{BA}\oplus_{\mathcal{I}}\mathcal{V}$, then the forward data-flow that starts from $\tilde{\mathbf{W}}$ is the same as in the full-rank training, as well as the backward data-flow that ends at $\tilde{\mathbf{W}}$. Therefore, the total compute of SLTrain should be $C_{\text{full-rank}}$ plus reconstructing $\tilde{\mathbf{W}}$, and its corresponding $2\times$ compute during backward, i.e.,
@@ Appendix C Detailed Memory Analysis
Methods Memory Re-Compute
Full-Rank $20nd+2n^{2}h$ N/A
Vanilla GCP ndnd $23nd^{2}+4n^{2}d$
CoLA $17.5nd+2n^{2}h+14nr$ N/A
CoLA-M $2nd+7nr$ $18.5ndr+4n^{2}d$
We analyze the memory and re-computation cost using the same notions as in Section 3.4 and denote $h$ as the number of attention heads. We further simplify the analysis under LLaMA architecture by uniformly assuming $d_{\text{ff}}\approx 2.5d$. We start with the activation memory of full-rank training:
@@
We summarize the results in Table 9.
Appendix D Training Configurations
@@ D.1 LLaMA Pre-Training
Loss QQP SST-2 MRPC COLA QNLI MNLI RTE STS-B GLUE Avg
[…]
CoLA – Gated MLP, Low-Rank $\sigma$ Only 1.257 91.2 92.3 90.6 54.1 91.7 84.3 74.2 89.7 83.5
CoLA – Original MLP, Preserve Full-Rank $\sigma$ 1.265 91.2 92.1 91.7 55.1 91.5 83.7 73.1 89.8 83.5
@@
[…]
The only caveat of adopting CoLA onto BERT is, we can’t remove the full-rank activation unless we modify its MLP structure, because the original BERT MLP has a two-layer structure, i.e., $\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x)}$. If we adopt CoLA while removing the full-rank activation, it becomes $\mathbf{B}_{2}\sigma\left(\mathbf{A}_{2}\mathbf{B}_{1}\sigma(\mathbf{A}_{1}\mathbf{x})\right)$, in which $\mathbf{A}_{2}$ and $\mathbf{B}_{1}$ are adjacent with no other operations in between, therefore $\mathbf{A}_{2}\mathbf{B}_{1}$ is mathematically equivalent to an $r$ to $r$ linear transformation. This could lead to a significant performance drop: we tried this setup at phase 1 (sequence length 128), resulting in a higher pre-training loss 1.579 vs 1.403. To avoid this setup, we naturally have two solutions: (1) use gated MLP, such as the ones in LLaMA/Mixtral/Qwen Models, then the full-rank activation can be safely removed; (2) use the original non-gated MLP, but preserve the full-rank activation on top of CoLA. In both solutions, a non-linear operation is placed between $\mathbf{A}_{2}$ and $\mathbf{B}_{1}$. We clarify that results shown in Section 5 are using solution (1).
To isolate the effect of CoLA from the structural change of the MLP layer, we also show results from solution (2) and compare with those from solution (1) side by side in Table 10. Interestingly, the average score for both solutions happen to be the same, with some variations in task-level performance. At similar performance, solution (1) yields lower FLOPs and memory cost, and is more aligned with CoLA’s design principal. Therefore, we promote solution (1).
@@ E.1 Ablation Study
CoLA w/ Only Low-Rank $\sigma$
– Reduced
35.41 25.90 20.50